OpenStudy (anonymous):

Part 1: Write an expression for the following real-world situation. If Mr. Jackson earns $1000 for every used car he sells and $2000 for every new car he sells, how much would he make if he sold "u" used cars and "n" new cars? Part 2: Write a real-world situation of your own and a VARIABLE expression to go with it.

OpenStudy (anonymous):

for this problem we'll start by making two equations, one that represents money made from used cars, and another that represents money made from new cars

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

so for every used car that Mr J sells he makes $1000, so if he sells two used cars he will make two times that amount, if he sells three he will make three TIMES that amount, so if he sells u amount of cars, he will make u times $1000

OpenStudy (anonymous):

so our first equation looks like this (number of used cars) times $1000 = amount of money made

OpenStudy (anonymous):

making sense so far?

OpenStudy (anonymous):

Yess

OpenStudy (anonymous):

good so our second equation is the same thing but with new cars, the number of new cars times $2000 equals the amount of money he makes (number of new cars) x $2000 = amount of money made from new cars

OpenStudy (anonymous):

so we have two equations \[u \times $1000 = m1\] and \[n \times $2000 = m2\]

OpenStudy (anonymous):

you see ive inserted variables, that make sense?

OpenStudy (anonymous):

Yess It Makes Sense

OpenStudy (anonymous):

ok so now we need to make a third equation, m1 + m2 = T (amount of money from used cars) + (amount from new cars) = (total amount of money)

OpenStudy (anonymous):

we then substitute the first two equations into the third to make (u x 1000) + (n x 2000) = T

OpenStudy (anonymous):

and that's your equation, does that make sense?

OpenStudy (anonymous):

Yea It Dose !

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

do you want help with the next part?